One canned juice drink is 15% orange juice, another is 10% orange juice. How many liters of each should be mixed together in order to get 5 liters that is 14% orange juice?
Answer by stanbon(75887)

One canned juice drink is 15% orange juice, another is 10% orange juice. How many liters of each should be mixed together in order to get 5 liters that is 14% orange juice?
0 solutions
----------------
Amount of 15% OJ = x liters ; amt of active ingredient = 0.15x liters
------------
Amount of 10% OJ="5-x" liters;amt of active ingredient=0.10(5-x)=0.5-0.1x lits
--------------------
Amount of 14% OJ=5 liters; amt of active ingredient = 0.14*5 = 0.7
-----------------------------
EQUATION:
active + active = active
0.15x + 0.5 - 0.10x = 0.7
15x + 50 - 10x = 70
5x=20
x=4 liters (amount of 15% OJ)
5-x = 1 liter (amount of 10% OJ)
===============
Cheers,
Stan H.